OpenStudy (cenaida):

Write the equation of the line with the given slope. 5/2 and y-intercept,-10

OpenStudy (anonymous):

do you know the slope intercept of a line?

OpenStudy (cenaida):

No

OpenStudy (anonymous):

\[y =mx +b\] where b is the slope and m is the y intercept

OpenStudy (cenaida):

Then is it D?

OpenStudy (anonymous):

i don't know.....what is D?

jimthompson5910 (jim_thompson5910):

jonnymiller, you have it mixed up m = slope b = y-intercept

OpenStudy (cenaida):

Y=-10x+(5/2)

OpenStudy (anonymous):

opps i was backwards! thanks don't know what i was thinking

jimthompson5910 (jim_thompson5910):

that's ok, silly typo

OpenStudy (anonymous):

m is slope and b is the y intercept!!!!

OpenStudy (cenaida):

A. Y=(5/2)x-10 B? Y= (5/2)x+10 C. Y=-10x+(5/2)

jimthompson5910 (jim_thompson5910):

if you choose B, you have the correct slope, but the wrong y-intercept

OpenStudy (cenaida):

Then A

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

correct